The records of an Italian shoe manufacturer show that 20% of the shoes are
defective. Assume that selection of shoes follow a Bernoulli trial.
a. If a manufacturer inspects 15 shoes randomly, what is the probability that at
most two shoes are defective?
b. If a manufacturer wants to find 6 non defective shoes, what is the probability
that 12 shoes have to be inspected?
c. Find the expected number of inspected shoes to find 6 non defective shoes.
1
Expert's answer
2020-04-03T14:13:10-0400

a) "P\{\xi\leq2\}=\sum_{s=0}^{2}C_{15}^{s}(0.2)^s(0.8)^{15-s}=0.398" (Bernoulli's formula where "n=15, k=2, p=0.2").

b) "P\{\eta=6\}=C_{12}^{6}(0.8)^6(0.2)^6=0.0155" (Bernoulli's formula where "n=12, k=6, p=0.8").

c) "M\zeta-?"

"M\zeta=np=n(0.8)=6\\
n=7.5\\
(\text{if }\zeta \in B(n,p)\text{ then } M\zeta=np)"